# Antipersistent energy current correlations in strong long-ranged Fermi-Pasta-Ulam-Tsingou type models

### 17 Jan 2024 v1

@@ Abstract
We study heat transfer in one-dimensional Fermi-Pasta-Ulam-Tsingou type systems with long-range (LR) interactions. The strength of the LR interaction between two lattice sites decays as a power $\sigma$ of the inverse of their distance. We focus on the strong LR regime ($0\leq\sigma\leq 1$) and show that the thermal transport behaviors are remarkably nuanced. Specifically, we observe that the antipersistent (negative) energy current correlation in this regime is intricately dependent on $\sigma$, displaying a nonmonotonic variation. Notably, a significant qualitative change occurs at $\sigma_{c}=1/2$, where with respect to other $\sigma$ values, the correlation shows a minimum negative value. The underlying mechanisms of this intriguing phenomenon is related to the two diverse space-time scaling properties of equilibrium heat correlations and the various scattering processes of phonons and discrete breathers.
@@ I Introduction
In studies of the LR-FPUT model, LR interactions can be added in two ways, i.e., LR quartic potential terms LRHeat-2017-1 ; LRHeat-ours-1 ; LRHeat-2021 ; LRHeat-ours-2 and a combination of both LR quadratic and quartic terms LRHeat-2018 ; LRHeat-2019 . These studies have revealed more interesting transport properties compared to the LR-rotor model. A notable discovery LRHeat-2017-1 is the behavior at $\sigma=2$, where thermal conductivity $\kappa$ shows an almost linear divergence with system size $L$ ($\kappa\sim L^{\alpha}$ with $\alpha=1$) and the temperature profile has a negligible slope. This ballistic-like transport hints at some yet unknown integrable limit of the model at $\sigma=2$ LRHeatreview-2 ; LRHeatreview-3 . However, this LR interacting system is a nonintegrable one. Our subsequent study LRHeat-ours-1 carefully considered the reservoir and boundary effects and observed a superdiffusive transport, instead, with a significantly high divergent exponent $\alpha\approx 0.7$, much larger than the $\alpha\approx 0.3-0.5$ typical of short-range interacting systems with nearest-neighbor (NN) couplings. This peculiar superdiffusive transport at $\sigma=2$ was further characterized by a slower decay in energy current correlation and linked to a weaker nonintegrability mechanism and the presence of a special type of traveling discrete breathers (DBs) in the system that enhance thermal transport LRHeat-ours-1 . Although this new type of superdiffusive energy transport was later doubted by Ref. LRHeat-2021 , the collective body of works LRHeat-2017-1 ; LRHeat-2018 ; LRHeat-ours-1 ; LRHeat-2021 do confirm the unique properties observed at $\sigma=2$.
@@
Beyond $\sigma=2$, it was initially posited that the LR-FPUT model exhibits superdiffusive transport across all other $\sigma$ values LRHeat-2017-1 . However, the parallel energy transport mechanism, proposed in LRHeat-2018 is likely taking its effects within $10<\sigma<$ 1 and a detailed scaling analysis of equilibrium correlation for $31<\sigma<$ 3 indicates a unique non-standard, non-local effective hydrodynamics, which differ from those observed in short-range interacting systems with NN couplings LRHeat-2019 . Therefore, it is quite strange that for $10<\sigma<$ 1, the LR-FPUT model demonstrates superdiffusive transport, while the LR-rotor model shows an insulating phase, indicative of subdiffusive energy transport. This disparity raises questions about the striking differences in transport properties between these two models LRHeat-2017-2 . To clarify this disparity, our recent study LRHeat-ours-2 has unequivocally corroborated the existence of subdiffusive energy transport in the mean-field limit of the LR-FPUT model, akin to that observed in the LR-rotor model.
@@
In view of these developments, the $2020$ study by Tamaki and Saito LRHeat-2020 aimed to derive a clear-cut analytical result that would complement the existing numerical findings. They conducted an analytical investigation of thermal transport properties in a linear LR-harmonic system, which involves stochastic momentum-exchange dynamics with both conserved and nonconserved momenta. Their idea is to calculate the equilibrium energy current correlation function, and then applying the Green-Kubo formula to determine thermal conductivity. However, their calculations were restricted to the weak LR regime, specifically for $\sigma\geq 2$ in momentum-conserving systems, and $\sigma\geq 3/2$ in momentum-nonconserving systems. For sufficiently strong LR systems, the divergence of the current correlation hampers their calculations. Despite this challenge, the work of Tamaki and Saito LRHeat-2020 has spurred further recent studies on systems without involving stochastic dynamics, such as the momentum-nonconserving LR-$\phi^{4}$ model LRHeat-2022-1 ; the momentum-conserving LR-harmonic systems in the mean-field limit LRHeat-2022-1 ; LRHeat-2022-2 and investigations into the strong LR regime beyond the mean-field case both using the nonequilibrium Green’s function formalism LRHeat-2023 .
@@
In this work we study numerically the thermal transport properties of LR-FPUT model, specifically focusing on the strong LR regime beyond the mean-field case ($0\leq\sigma\leq 1$)—a domain not yet addressed by current analytical methods. Instead of examining the system size-dependent thermal conductivity as in previous studies LRHeat-2016 ; LRHeat-2017-1 ; LRHeat-2018 , our study concentrates on calculating the equilibrium energy current correlation function. Since in Ref. LRHeat-ours-2 we have observed the subdiffusive energy transport and its antipersistent energy current correlation in the LR-FPUT model in the mean-field case, and our results are distinct from the superdiffusive transport observed in LRHeat-2017-1 , in the present work we also wish to explore further how this antipersistent energy current correlation should change with the variation of $\sigma$.
@@ II Model and method
In this Hamiltonian, $x_{i}$ and $p_{i}$ are two canonically conjugated variables with $i$ the index of the particle; all other relevant quantities like the particle’s mass and the lattice constant are dimensionless and set to be unity. The periodic boundary conditions used make our system like a ring which ensures that our study does not suffer from any uncertain effects caused by the open boundary conditions. The interparticle interaction has two separate terms as our previous works LRHeat-ours-1 ; LRHeat-ours-2 . The NN couplings are added to the quadratic potential term and the LR interactions are included in the quartic potential term. In the LR potential term, $r^{-\sigma}$ represents the interaction strength of the $i$th particle with its $r$th neighborings with $\sigma$ the range value of LR exponent. Such a setup of the LR-FPUT model has been found to showing more interesting dynamic phenomena than that by a combination of both the LR quadratic and quartic potential terms JianWang2022 ; JianWang2022cite .
@@
$J_{\rm tot}=\sum_{i=1}^{N}p_{i}\left[(x_{i+1}-x_{i})+\sum_{r=1}^{\frac{N}{2}-1}(x_{i+r}-x_{i})^{3}\right]$ (4)
@@ III Range-dependent energy current correlations
Figure 1: The equilibrium heat current autocorrelation $C_{JJ}(t)$ versus $t$ for several $\sigma$ within the LR strong regime (with $0\leq\sigma\leq 1$). Here for a better visualization, the values of the correlations have been rescaled between 00 and $1$ by divided by $C_{JJ}(0)$.
@@ IV Space-time scaling properties of equilibrium heat correlations
Figure 2: The equilibrium spatiotemproal correlation function of heat $\rho_{Q}(m,t)$ for several $0\leq\sigma\leq 1$ for three typical long times: $t=500$ (dotted), $t=1000$ (dashed), and $t=2000$ (solid).
@@
Figure 2 depicts $\rho_{Q}(m,t)$ for three long times for several $0\leq\sigma\leq 1$. First, for all of $\sigma$ considered, it shows a relatively slow spread of heat [all $\rho_{Q}(m,t)$ for a time $t=2000$ only cover few lattice sites, which are much smaller than the system size], indicating the possible mechanism of subdiffusive thermal transport as shown in Fig. 1. Again, this is distinct from the superdiffusive transport observed in Ref. LRHeat-2017-1 . Another information is that with the increase of $\sigma$, especially for $\sigma=0.8$ and $\sigma=1$ [see Figs. 2(e, f)], $\rho_{Q}(m,t)$ appears to be delocalized, indicating the absence of subdiffusive energy transport. This is consistent with the observation of energy current correlations in Fig. 1.
@@
We now try to understand why $\sigma_{c}=0.5$ is a special point. We will not study the mean squared displacement of $\rho_{Q}(m,t)$ versus $t$ here because of the relatively large fluctuations of $\rho_{Q}(m,t)$ for a large distance $m$. Instead, we will focus on the decay of the central peak $\rho_{Q}(0,t)$ of $\rho_{Q}(m,t)$. Such a decrease of $\rho_{Q}(0,t)$ versus $t$ has been used to infer the long-time asymptotic behavior of heat spread for the short-range systems with NN couplings DX2018 ; DX2017-1 ; DX2017-2 and for the LR-FPUT models LRHeat-2017-1 ; LRHeat-ours-1 ; LRHeat-2021 in the regimes of ballistic, superdiffusive, and diffusive transport. In those cases, if $\rho_{Q}(0,t)\sim t^{-\gamma}$, then based on the scaling formula of the single particle Lévy walk model Report2015 we have
@@
However, for subdiffusive thermal transport considered here, one does not have such a scaling law $\rho_{Q}(0,t)\sim t^{-\gamma}$. Fortunately, in Ref LRHeat-ours-2 we for the first time revealed that in the LR-FPUT model of the mean-field case ($\sigma=0$) following subdiffusive thermal transport, an exponential decay law $\rho_{Q}(0,t)\sim\exp(-\frac{t}{\tau})$ is available. Therefore, bearing the results of Figs. 1 and 2 in mind, one may expect a significant change between these two distinct laws for $\rho_{Q}(0,t)$ versus $t$ as $\sigma$ increases.
To confirm this expectation, in Fig. 3 the results of $\rho_{Q}(0,t)$ versus $t$ are plotted in the following two ways: (a) a single log plot and (b) a log-log plot. As can be seen, there is indeed a crossover between the two distinct laws, which characterize different transport regimes. More-interestingly, from Fig. 3 one can easily capture that $\sigma_{c}=0.5$ is a special point separating the laws of $\rho_{Q}(0,t)\sim\exp(-\frac{t}{\tau})$ and $\rho_{Q}(0,t)\sim t^{-\gamma}$. This finding supports the peculiarity of $\sigma_{c}=0.5$ as indicated in the antipersistent energy current correlations in Fig. 1.
V Scattering dynamics of phonons and DBs

### 22 Mar 2024 v3

@@ Abstract
We study heat transfer in one-dimensional Fermi-Pasta-Ulam-Tsingou type systems with long-range (LR) interactions. The strength of the LR interaction between two lattice sites decays as a power $\sigma$ of the inverse of their distance. We focus on the strong LR regime ($0\leq\sigma\leq 1$) and show that the thermal transport behaviors are remarkably nuanced. Specifically, we observe that the antipersistent (negative) energy current correlation in this regime is intricately dependent on $\sigma$, displaying a nonmonotonic variation. Notably, a significant qualitative change occurs at $\sigma_{c}=0.5$, where with respect to other $\sigma$ values, the correlation shows a minimum negative value. Furthermore, our findings also demonstrate that within the long-time range considered, these antipersistent correlations will eventually vanish for certain $\sigma>0.5$. The underlying mechanisms behind these intriguing phenomena are related to the crossover of two diverse space-time scaling properties of equilibrium heat correlations and the various scattering processes of phonons and discrete breathers.
@@ I Introduction
In studies of the LR-FPUT model, LR interactions can be added in two ways, i.e., LR quartic potential terms LRHeat-2017-1 ; LRHeat-ours-1 ; LRHeat-2021 ; LRHeat-ours-2 and a combination of both LR quadratic and quartic terms LRHeat-2018 ; LRHeat-2019 . These studies have revealed more interesting transport properties compared to the LR-rotor model. A notable discovery LRHeat-2017-1 is the behavior at $\sigma=2$, where thermal conductivity $\kappa$ shows an almost linear divergence with system size $L$ ($\kappa\sim L^{\alpha}$ with $\alpha=1$) and the temperature profile has a negligible slope. This ballistic-like transport hints at some yet unknown integrable limit of the model at $\sigma=2$ LRHeatreview-2 ; LRHeatreview-3 . However, this LR interacting system is a nonintegrable one. Our subsequent study LRHeat-ours-1 carefully considered the reservoir and boundary effects and observed a superdiffusive transport, instead, with a significantly high divergent exponent $\alpha\approx 0.7$, much larger than the $\alpha\approx 0.3-0.5$ typical of short-range interacting systems with nearest-neighbor (NN) couplings only. This peculiar superdiffusive transport at $\sigma=2$ was further characterized by a slower decay in energy current correlation and linked to a weaker nonintegrability mechanism and the presence of a special type of traveling discrete breathers (DBs) in the system that enhance thermal transport LRHeat-ours-1 . Although this new type of superdiffusive energy transport was later doubted by Ref. LRHeat-2021 , the collective body of works LRHeat-2017-1 ; LRHeat-2018 ; LRHeat-ours-1 ; LRHeat-2021 do confirm the unique properties observed at $\sigma=2$.
@@
Beyond $\sigma=2$, it was initially posited that the LR-FPUT model exhibits superdiffusive transport across all other $\sigma$ values LRHeat-2017-1 . However, the parallel energy transport mechanism, proposed in LRHeat-2018 is likely taking its effects within $10<\sigma<$ 1 and a detailed scaling analysis of equilibrium correlation for $31<\sigma<$ 3 indicates a unique non-standard, non-local effective hydrodynamics, which differ from those observed in short-range interacting systems with NN couplings only LRHeat-2019 . Therefore, it is quite strange that for $10<\sigma<$ 1, the LR-FPUT model demonstrates superdiffusive transport, while the LR-rotor model shows an insulating phase, indicative of subdiffusive energy transport. This disparity raises questions about the striking differences in transport properties between these two models LRHeat-2017-2 . To clarify this disparity, our recent study LRHeat-ours-2 has unequivocally corroborated the existence of subdiffusive energy transport in the mean-field limit of the LR-FPUT model, akin to that observed in the LR-rotor model.
@@
In view of these developments, the $2020$ study by Tamaki and Saito LRHeat-2020 aimed to derive a clear-cut analytical result that would complement the existing numerical findings. They conducted an analytical investigation of thermal transport properties in a linear LR-harmonic system, which involves stochastic momentum-exchange dynamics with both conserved and nonconserved momenta. Their idea is to calculate the equilibrium energy current correlation function, and then applying the Green-Kubo formula to determine thermal conductivity. However, their calculations were restricted to the weak LR regime, specifically for $\sigma\geq 2$ in momentum-conserving systems, and $\sigma\geq 3/2$ in momentum-nonconserving systems. For sufficiently strong LR systems, the divergence of the current correlation hampers their calculations. Despite this challenge, the work of Tamaki and Saito LRHeat-2020 has spurred further recent studies on systems without involving stochastic dynamics, such as the momentum-nonconserving LR-$\phi^{4}$ model LRHeat-2022-1 ; the momentum-conserving LR-harmonic systems in the mean-field limit LRHeat-2022-2 ; LRHeat-2022-3 and investigations into the strong LR regime beyond the mean-field case both using the nonequilibrium Green’s function formalism LRHeat-2023 .
@@
In this work we study numerically the thermal transport properties of LR-FPUT model, specifically focusing on the strong LR regime beyond the mean-field case ($0\leq\sigma\leq 1$)—a domain not yet addressed by current analytical methods. Instead of examining the system size-dependent thermal conductivity as in previous studies LRHeat-2016 ; LRHeat-2017-1 ; LRHeat-2018 , our study concentrates on calculating the equilibrium energy current correlation function. Since in Ref. LRHeat-ours-2 we have observed the subdiffusive energy transport and its antipersistent energy current correlation in the LR-FPUT model in the mean-field case, and our results are distinct from the superdiffusive transport observed in LRHeat-2017-1 , in the present work we also wish to explore further how this antipersistent energy current correlation would change with the variation of $\sigma$.
@@ II Model and method
In this Hamiltonian, $x_{i}$ and $p_{i}$ are two canonically conjugated variables with $i$ the index of the particle; all other relevant quantities like the particle’s mass and the lattice constant are dimensionless and set to be unity. The periodic boundary conditions used make our system like a ring which ensures that our study does not suffer from any uncertain effects caused by the open boundary conditions. The interparticle interaction has two separate terms as our previous works LRHeat-ours-1 ; LRHeat-ours-2 . The NN couplings are added to the quadratic potential term and the LR interactions are included in the quartic potential term. In the LR potential term, $r^{-\sigma}$ represents the interaction strength of the $i$th particle with its $r$th neighborings with $\sigma$ the range value of LR exponent. Such a setup of the LR-FPUT model including the LR quartic potential term only has been found to showing more interesting dynamic phenomena than that by a combination of both the LR quadratic and quartic potential terms JianWang2022 ; JianWang2022cite , and therefore, in the present work we will focus on this case only.
@@
$J_{\rm tot}=\sum_{i=1}^{N}p_{i}\left[(x_{i+1}-x_{i})+\sum_{r=1}^{\frac{N}{2}-1}\frac{(x_{i+r}-x_{i})^{3}}{r^{\sigma}}\right]$ (4)
@@ III Range-dependent energy current correlations
Figure 1: The equilibrium heat current autocorrelation $C_{JJ}(t)$ versus $t$ for several $\sigma$ within the LR strong regime (with $0\leq\sigma\leq 1$). Here for a better visualization, the correlation values have been rescaled between 00 and $1$ by dividing them with $C_{JJ}(0)$.
@@ IV Space-time scaling properties of equilibrium heat correlations
Figure 2: The equilibrium spatiotemproal correlation function of heat $\rho_{Q}(m,t)$ for several $0\leq\sigma\leq 1$ for three typical long times: $t=1000$ (dotted), $t=2000$ (dashed), and $t=4000$ (solid).
@@
Figure 2 depicts $\rho_{Q}(m,t)$ for three long times for several $0\leq\sigma\leq 1$. First, for all of $\sigma$ considered, it shows a relatively slow spread of heat [all $\rho_{Q}(m,t)$ for a time $t=4000$ only cover few lattice sites, which are much smaller than the system size], indicating the possible mechanism of subdiffusive thermal transport as shown in Fig. 1. Again, this is distinct from the superdiffusive transport observed in Ref. LRHeat-2017-1 . Another information is that with the increase of $\sigma$, especially for $\sigma=0.8$ and $\sigma=1$ [see Figs. 2(e, f)], $\rho_{Q}(m,t)$ appears to be delocalized, indicating the absence of subdiffusive energy transport. This is consistent with the observation of energy current correlations in Fig. 1.
@@
We now try to understand why $\sigma_{c}=0.5$ is a special point. We will not study the mean squared displacement of $\rho_{Q}(m,t)$ versus $t$ here because of the relatively large fluctuations of $\rho_{Q}(m,t)$ for a large distance $m$. Instead, we will focus on the decay of the central peak $\rho_{Q}(0,t)$ of $\rho_{Q}(m,t)$. Such a decrease of $\rho_{Q}(0,t)$ versus $t$ has been used to infer the long-time asymptotic behavior of heat spread for the short-range systems with NN couplings only DX2018 ; DX2017-1 ; DX2017-2 and for the LR-FPUT models LRHeat-2017-1 ; LRHeat-ours-1 ; LRHeat-2021 in the regimes of ballistic, superdiffusive, and diffusive transport. In those cases, if $\rho_{Q}(0,t)\sim t^{-1/\gamma}$, based on the scaling formula of the single particle Lévy walk model Report2015 we have
@@
Then, $\gamma=1$, $21<\gamma<$ 2, and $\gamma=2$ correspond to ballistic, superdiffusive, and diffusive transport, respectively, where one finds $1/2\leq 1/\gamma\leq 1$. However, for subdiffusive thermal transport considered here, one does not have such a scaling law of $\rho_{Q}(0,t)\sim t^{-1/\gamma}$. Fortunately, in Ref LRHeat-ours-2 we for the first time revealed that in the LR-FPUT model of the mean-field case ($\sigma=0$) following subdiffusive thermal transport, a two-stage exponential decay law $\rho_{Q}(0,t)\sim\exp(-\frac{t}{\tau})$ is available. Therefore, bearing the results of Figs. 1 and 2 in mind, one may conjecture a crossover between these two distinct laws for $\rho_{Q}(0,t)$ versus $t$ as $\sigma$ increases.
To provide evidence for our conjecture, Fig. 3 presents the results of $\rho_{Q}(0,t)$ versus $t$ in two distinct ways: (a) a single-log plot and (b) a log-log plot. As can be seen, both results of $\sigma=0$ and $\sigma=1$ for long times can be well fitted by a line. This indicates that there is indeed a crossover between the two distinct laws, which characterizes different transport regimes. Furthermore, since $\sigma_{c}=0.5$ lies between the extremes of $\sigma=0$ and $\sigma=1$, this finding provides preliminary evidence supporting the speciality of $\sigma_{c}=0.5$, as indicated by the antipersistent energy current correlations depicted in Fig. 1.
Figure 4: $\rho_{Q}(0,t)$ versus $t$ in a single-log plot showing the detailed two-stage exponential decay for (a) $\sigma=0$, (b) $\sigma=0.25$, (c) $\sigma=0.4$, and (d) $\sigma=0.5$.
From Fig. 1 the antipersistent energy current correlations mainly occur in a relatively short time, it is thus necessary to examine more detailed variation of the scaling of $\rho_{Q}(0,t)$ in a short time. Based on the current data with a time up to $t=4000$, we have tried both plots of single-log and log-log. We find that the single-log plot is helpful to highlight the short-time behavior. Therefore, in the following we will first present the details of the two-stage exponential decay behavior of $\rho_{Q}(0,t)$ for $\sigma\in[0,0.5]$, then focus on the results on the vicinity of $\sigma_{c}=0.5$, and finally provide the power-law decay of $\rho_{Q}(0,t)$ for $\sigma\in(0.5,1]$. Figure 4 depicts the two-stage exponential decay behavior for four typical $\sigma$ values within $[0,0.5]$. As to this figure, we first recall the known results of $\sigma=0$. Indeed, similar to the observations for the equilibrium spatiotemporal correlations of total energy in LRHeat-ours-2 , we see a perfect two-stage exponential decay characterized by two distinct timescales: $\tau_{1}\simeq 1000$ for short times and $\tau_{2}\simeq 1445$ for long times [see Fig. 4(a)]. In particular, this perfect two-stage exponential decay has been conjectured to related to the two different energy relaxation induced by the system’s surviving DBs, before and after phonon-DBs scattering, respectively. In this conjecture, due to the significantly stronger strength of DBs relative to phonons in the mean-field scenario, phonon-DB scattering acts as a transient process which leads to a rapid switching between the two exponential decays as mentioned in LRHeat-ours-2 and also shown in Fig. 4(a). This may be the reason why we observe a relatively small value of $|C_{JJ}(t)|$ at its negative minimum for $\sigma=0$ if compared to other $\sigma\in[0,0.5]$ [see Fig.1].
If we acknowledge the reasonability of the above conjecture, let us next discuss the results depicted in Figs. 4(b)-(d). The key distinction if compared to $\sigma=0$ is that the switching between the two exponential decays now becomes progressively slower. This may imply a prolonged persistence of phonon-DBs scattering, resulting in an increased reflection of heat flux and consequently leading to a larger value of $|C_{JJ}(t)|$ at its negative minimum if compared to $\sigma=0$. Interestingly, with the range $[0,0.5]$ considered, $\sigma=0.5$ corresponds to the slowest switching [see Fig. 4(d)]. This fact may explain why the peculiarity of $\sigma_{c}=0.5$ occurs at least for $0\leq\sigma\leq 0.5$. We also find that from Fig. 4(d) wherein $\sigma=0.5$ the time period associated with the initial-stage exponential decay is relatively short. This suggests strong scattering between phonons and DBs, which hinders the formation of an exponential decay during this first stage. On the other hand, the second-stage exponential decay seems still available, which may indicate that DBs would likely serve as predominant energy carriers after the phonon-DBs scattering as the foundation of exponential decay of $\rho_{Q}(0,t)$ should be attributed to DBs.
Figure 5: $\rho_{Q}(0,t)$ versus $t$ in a single-log plot showing the detailed variation around $\sigma=0.5$.
We now turn to the results on the vicinity of $\sigma_{c}=0.5$, as shown by a single-log plot of $\rho_{Q}(0,t)$ versus $t$ in Fig. 5. We note again that a log-log plot would be more helpful to examine the long-time behavior, while from Fig. 1 the antipersistent energy current correlations only happens at short times, we therefore choose to focus on a single-log plot here. From Fig. 5 one may identify the special difference of $\sigma=0.5$, if compared to the results obtained for $\sigma=0.4$ and $\sigma=0.6$. Firstly, based on the current data, while all values of $\rho_{Q}(0,t)$ for $\sigma=0.4$ for different times are always larger than those for $\sigma=0.5$; for $\sigma=0.6$, initially we see a smaller $\rho_{Q}(0,t)$, but in a long time it will be reversed and such reversion becomes even more prominently for $\sigma=0.8$. Secondly, beyond $\sigma=0.5$, the inital-stage exponential decay exhibited by $\rho_{Q}(0,t)$ seems almost disrupted and no longer observable (see the result of $\sigma=0.8$). In particular, there exhibits a long tail for $\sigma=0.8$, instead. Based on this long tail, we further conjecture that unlike $0\leq\sigma\leq 0.5$, for $\sigma>0.5$ phonons may dominate over DBs after undergoing phonon-DBs scattering due to their stronger strength. If this is the case, the relaxation process induced by phonons (perturbed by DBs) would follow a distinct behavior if compared to solely considering DBs. Therefore, even though an antipersistent $C_{JJ}(t)$ can still be observed in certain cases within $\sigma>0.5$ (as evidenced from Fig. 1), as $\sigma$ increases further, the long tail of $\rho_{Q}(0,t)$ for $\sigma>0.5$ would become more evident (see the result of $\sigma=0.8$). Eventually, $\sigma_{c}=0.5$ may serve as a change point characterizing whether DBs or phonons act as primary surviving energy carriers after phonon-DBs scattering. This may give a full story for why $\sigma_{c}=0.5$ exhibits peculiar behavior in the whole range of $\sigma\in[0,1]$, and our findings depicted in Figs. 4 and 5 seem to support this conjecture.
Figure 6: $\rho_{Q}(0,t)$ versus $t$ in a log-log plot showing the detailed power-law decay in long times for (a) $\sigma=0.6$, (b) $\sigma=0.7$, (c) $\sigma=0.8$, and (d) $\sigma=1$.
To capture the observed long tail exhibited in $\rho_{Q}(0,t)$ for $\sigma>0.5$, similar to $\sigma=0.8$, Fig. 6 presents a log-log plot of $\rho_{Q}(0,t)$ versus $t$, providing detailed information for several typical $\sigma$ within $\sigma>0.5$. Since at present only a long time up to $t=4000$ is available, we should admit that to explore the long-time asymptotic behaviors for a $\sigma$ value close to $\sigma_{c}=0.5$ is still challenging, but the results of Fig. 6 do provide further insights. As can be seen in Fig. 6, based on the current data, while the best fittings of $1/\gamma$ for $\sigma=0.6$ and $\sigma=0.7$ exceed 1 [see Figs. 6(a) and (b)], only for $\sigma\geq 0.8$ [see Fig. 6(c)] does $1/\gamma$ approximately equal to $1/\gamma\simeq 1$, which now falls within the theoretical prediction range of $1/2\leq 1/\gamma\leq 1$ according to the Lévy walk model Report2015 . Furthermore, as $\sigma$ increases further, one obtains an exponent close to the theoretical prediction of $\sigma=5/3$ for short-range FPUT systems with NN coupling only NNHeatreview-1 ; NNHeatreview-2 ; NNHeatreview-3 , with $1/\gamma\simeq 0.61$. Such findings are consistent with the transition from localization to delocalization as shown in Fig. 2. Additionally, they indicate that apart from considering $\gamma_{c}=0.5$, there is a $\gamma$ ($\gamma>0.5$) value, probably being $\sigma\simeq 0.8$ (within the time range considered), above which, the long tail will emerge and the antipersistent energy current correlations will be absent (see also Fig. 1).
V Absence of antipersistent energy current correlation